# GMIMS-DRAGONS: A Faraday Depth Survey of the Northern Sky Covering 350 to 1030 MHz

### v1

@@ 2.2 The 350–1030 MHz Receiver
In the frequency band of this survey, commercially available devices were able to provide adequate receiver sensitivity and bandwidth. Low-noise amplifiers were followed by conventional amplification and band-equalizing stages placed on the feed platform. The receiver passband was defined by the cut-off frequency of the feed at ${\sim}350$ MHz, and by a low-pass filter, with 3 dB frequency ${\sim}1030$ MHz, placed at the end of the feed-platform amplifier chain. The amplified outputs were transported to the control building by an RF-over-fiber (RFoF) device. Signals were further amplified to 20-20 dBm, the specified input level to the analog-to-digital converters. The calibration signal was provided by a solid-state noise source heated to a constant temperature near $30\arcdeg$ C. [^2] Outside nighttime temperatures did not exceed $30\arcdegC$ during the survey, and it is simpler to only heat the noise source rather than also implementing cooling in warmer weather.
2.3 The Digital Spectrometer
@@
Since the key data product from this survey is a Faraday depth cube, the requirements for the spectral channelization are primarily driven by bandwidth depolarization effects (Raja, 2014; Schnitzeler & Lee, 2015; Pratley & Johnston-Hollitt, 2020), which could, in principle, be met by a spectrometer that analyzes incoming signals to a frequency resolution of ${\sim}1$ MHz (see Section 6.2 for details). Faraday synthesis on such data produces a reliable Faraday depth cube (and we ultimately bin the channels to 1 MHz in the map-making stage). As well, digitization to four bits would satisfactorily limit any additional noise from the quantization process. However, in the real world, our observing frequencies carry heavy communications traffic, and we need much finer digitization in order to have the dynamic range to detect the astronomical signal in the presence of the much stronger radio frequency interference (RFI) signals. Channels and samples spoiled by RFI must be excised from the data, and to minimize data loss it is useful to have a sampling that is much finer than required by the demands of Faraday synthesis. Fortunately, modern digital electronics permits the design of spectrometers with many bits and many narrow frequency channels.
@@
Two of the 137.5 MHz-wide bands lie below the cutoff frequency of the feed and therefore do not contain useful data. The number of spectrometer channels used is about 16,300. Inter-channel isolation (between 41.6 kHz channels) is 70 dB, and passband ripple is 0.1 dB. Aliasing at the 1030 MHz roll-off of the low-pass filter is ${\sim}40$ dB. The four correlated outputs produced by the spectrometer are the total power from the left- and right-circular channels, RR* and LL*, and the linear polarization products, RL* and LR*.
2.4 Telescope Control System
Survey observations were made as azimuth scans at two fixed elevations (Section 3.1). With its azimuth-elevation configuration, the DRAO-15 telescope has the capability to scan quickly in azimuth (up to 180 deg/minute) at a fixed elevation. The azimuth axis uses two drives to prevent backlash. [^3] Discrepancy in the motion due to the clearance between the gear teeth. The elevation drive has a maximum speed of 60 deg/minute, allowing the telescope to change between the two survey elevations ($49\arcdeg$ and $20\arcdeg$; Section 3.1) in less than 4 minutes (slow-down and settling time included), permitting efficient scheduling of scans at the two elevations used for the survey within an observing night.
The low-level control software was written in C++, while the observing schedule for the survey (including azimuth scans and calibration scans) consisted of a set of python scripts. For the cross-scans for the pointing model (Section 2.5) and the raster scans of bright calibrators (Section 3.2.2) python scripts provided the telescope with a series of pointings in right ascension (RA) and declination, updated every 0.5 seconds, tracing out the pattern of the scan by commanding the telescope to ‘chase’ the changing target pointing.
Figure 4: Right ascension (top) and declination (bottom) offsets between the expected and observed pointing positions as a function of azimuth. These measurements are derived from Gaussian fits to the calibration raster scans of Cyg A, Cas A, Tau A, and Vir A (Section 3.2.2) produced over the course of the survey.
2.5 Pointing Model
@@
Each observation consisted of a ‘cross-scan’: one $8\arcdeg$-long scan in RA and one $8\arcdeg$-long scan in declination across the expected source position. Peak positions of one-dimensional Gaussian functions fitted to each scan in the LL* correlation product at frequencies above 900 MHz determined the measured positions. The six model parameters fitted to the data with TPoint were: azimuth offset, elevation offset, non-perpendicularity between the azimuth and elevation axes, elevation-dependent azimuth shift, azimuth-axis East-West misalignment and azimuth-axis North-South misalignment. After applying the corrections we re-observed a smaller subset of sources in order to assess the new pointing accuracy. The combined corrections produced a substantial improvement, reducing the systematic offsets between the expected and observed source positions by $\sim 80\%$ in both elevation and azimuth, and reducing the rms errors by $6\%$ in elevation and by $50\%$ in azimuth. The overall residual uncertainty in pointing is $\sim 5\%$ of the beamwidth, making this a sufficiently accurate correction for mapping large-scale structures.
@@
We applied the pointing model at the beginning of the survey and no further corrections were made throughout. However the daily calibration raster scans using the four brightest sources (Cyg A, Cas A, Tau A, and Vir A; Section 3.2.2) provided continuous checks on the pointing accuracy. We show the RA and declination offsets in Figure 3 as a function of frequency based on an averaged set of Cyg A raster scans, and in Figure 4 at 820 MHz as a function of telescope azimuth from the full set of raster scans. While there is a slight residual pointing accuracy dependence on both frequency and azimuth, the pointing accuracy remains within 5% of the beamwidth over the course of the survey, indicating sufficient pointing stability.
2.6 System and Noise Source Temperature
@@
Separate from the calibration measurements and procedures that we ultimately applied to the survey data (described in Section 3.2), we measured the noise temperature of the receiver and the noise equivalent of the injected calibration signal prior to installing the receiver on the telescope. This was accomplished using the DRAO Hot-Cold Test Facility (HCTF; Hovey et al., 2018). The HCTF is a large metal funnel, with its wide mouth open at the top. The floor of the funnel is a flat metal plate, 1 m square, and the top opening is 3 m square, 3 m above the floor. The feed and receiver were placed in it, pointing upwards. Noise measurements were put on an absolute scale by exposing the feed and receiver to two terminations at known temperatures. A hot termination consists of a sliding roof lined with microwave absorber, and a cold termination was provided by sliding the roof open, exposing feed and receiver to the sky. We measured the physical temperature of the hot absorber directly, and derived the brightness temperature of the sky from the PyGDSM sky model (Price, 2016; Zheng et al., 2017). Measurements were made with laboratory spectrum analyzers with frequency resolution of 3 MHz.
The receiver temperature and calibration noise-source temperature are shown in Figure 5. The receiver temperature varies smoothly between $\sim 80$ K at the high and low ends of the band and $\sim 55$ K at 600 MHz. The calibration noise-source temperature fluctuates between $\sim 15$ and $\sim 80$ K. The large-scale, sinusoidal fluctuations in the noise-source temperature are likely caused by reflections between the noise-source coupler (Section 3.2.1) and the metal plate to which it was mounted. We estimate the accuracy of these measurements as ${\pm}3$ K above 500 MHz. Accuracy below 500 MHz is limited by the size of the ${\sim}45$  cm-long pyramidal cones of the absorber that line the sliding roof. To investigate the effects of aliasing into the band we repeated the HCTF measurements out to 4 GHz, revealing that a negligible 0.05 to 0.3 K would be added to the system temperature by aliasing.
We calculate the expected noise in the final 1 MHz channel widths used for published data cubes using the radiometer equation applied to the receiver temperature values in Figure 5 and approximate sky brightness temperatures from the PyGDSM sky model (Price, 2016; Zheng et al., 2017). An integration time of $2\times 0.6=1.2$ s [^5] The effective integration time is twice the spectrometer integration time since each pointing is sampled twice in the scan strategy (Section 3.1). and 1 MHz channel-width yields an expected noise between 55 and 100 mK.
@@
Figure 5: The receiver noise temperature in kelvins (top) and the kelvin equivalent of the calibration noise source (bottom) as determined from the HCTF measurements. Results are shown for both right- and left-hand circular polarization (RCP; black and LCP; grey).
@@
2.7 Aperture Efficiency
As described in Section 4.8, we achieved conversion to brightness temperature (K) by calculating the main beam solid angle, as a function of frequency, from the measured beamwidths, assuming that the main beam is Gaussian. Nevertheless, we carried out an absolute calibration of the telescope, with the aim of improving our understanding of the instrument. We compared the antenna temperature produced by Cyg A with absolute noise standards, resistors at the temperature of liquid nitrogen, at ambient temperature, and in a temperature controlled oven at approximately $100^{\circ}$ C. We made this measurement using the RR* beam only. In the top panel of Figure 6 we show the measured aperture efficiency, ${\eta}_{A}$, compared with that computed using the General Reflector antenna Analysis Software Package by TICRA (GRASP) [^6] https://www.ticra.com/software/grasp/ electromagnetic simulator. The excellent agreement, generally within 1%, gives us great confidence in the GRASP computations of the radiation pattern of the antenna, which will be useful for future improvements to the ground emission calculations described in Section 4.4.
3 Observations
@@ 3.1 Observing Strategy
We used the observing technique developed by Carretti et al. (2019), scanning in azimuth at fixed elevation. The timing was selected such that scans adjacent to each other on the sky provided close to uniform spatial coverage, and the length of the scans was chosen such that scans would also cross one another, allowing for the possibility of basket-weaving to reconcile their base levels (see Section 4.6). A scan consists of all the data collected for one sweep in azimuth with the telescope moving either clockwise or counterclockwise over $360^{\circ}$ at $20^{\circ}$ per minute, starting and ending at azimuth ${a}={0^{\circ}}$ (North). Scans were made at two elevations, ${e}={49.3^{\circ}}$, the elevation of the North Celestial Pole (NCP) at DRAO, and at ${e}={20^{\circ}}$, just above the lower elevation limit for the telescope. We denote these as ${e}_{\text{high}}$ and ${e}_{\text{low}}$ scans respectively. The ${e}_{\text{high}}$ scans covered the declination range $11\arcdeg\leq\delta\leq 90\arcdeg$ and ${e}_{\text{low}}$ scans covered $20\arcdeg\leq\delta\leq 60\arcdeg$ (Figure 7). In the map-making stage we split each scan into ‘East’ ($0\arcdeg\leq{a}\leq 180\arcdeg$) and ‘West’ ($180\arcdeg\leq{a}\leq\ 360\arcdeg$) halves (Figure 8) which produced the intersecting scan patterns (Figure 9). We produced separate maps from the ${e}_{\text{high}}$ and ${e}_{\text{low}}$ datasets. We then combined these using a weighted average in the overlap region $11\arcdeg\leq\delta\leq\ 60\arcdeg$, with the ${e}_{\text{high}}$ (${e}_{\text{low}}$) map weights changing linearly from 0 to 1 (1 to 0) across $\delta=11\arcdeg$ to $\delta=60\arcdeg$, thereby producing the final maps covering $20\arcdeg\leq\delta\leq 90\arcdeg$.
@@
Scans started at sidereal times evenly spaced by an interval equivalent to 20 arcminutes on the celestial sphere. This guaranteed full Nyquist sampling of the sky at the highest survey frequency (1030 MHz) where the HPBW is $\sim 1\arcdeg$. Examples of the spacing between scans for different declination ranges in the ${e}_{\text{high}}$ and ${e}_{\text{low}}$ scans are shown in Figure 9. Apart from a few instances of missed scans, scan separations are typically at most 20′. The survey was completed in three phases over the course of seven months: phase 1 (P1; June–July, 2022), phase 2 (P2; August–October 2022) and phase 3 (P3; December 2022–January 2023). Observations were made at night in order to avoid contamination from solar emission received through the sidelobes. Although the offset reflector has low sidelobes, the Sun in the sidelobes can contribute spurious polarized signals as bright as the brightest Galactic emission, which we observed when running a short test version of the survey that included daytime data.
@@
In the survey planning process we generated a ‘library’ of scans, following Wolleben et al. (2021), with each scan having a unique 4-digit scan ID. This consisted of 1440 ${e}_{\text{low}}$ and 1440 ${e}_{\text{high}}$ scans. For each phase of the survey we generated a daily scan schedule, drawing from the scans with sidereal start times occurring after sunset and sidereal end times occurring before sunrise. These schedules were generated with two objectives: (i) to achieve survey efficiency, and (ii) to add some randomness to the scan sequence. These objectives were, obviously, in conflict, so we adopted a reasonable compromise, aiming to avoid fixed patterns in the observing sequence. Sets of both ${e}_{\text{high}}$ and ${e}_{\text{low}}$ scans were completed on each night. After each survey night, if a scan was deemed unsuccessful (e.g. a result of excessive RFI or equipment malfunction) the scan ID was returned to the master scan library to be rescheduled on a subsequent night. In total, 1401 of the 1440 ${e}_{\text{high}}$ and 1406 of the 1440 ${e}_{\text{low}}$ scans were completed for inclusion in the mapping stage. We made efforts to reschedule failed scans, but were in some cases limited by lack of available nighttime coverage of the required sidereal observing times. In the processing steps we also discarded a small number of scans from individual frequency channels based on an rms filtering threshold (Section 4.3).
3.2 Calibration Observations

### v2

@@ II.2 The 350–1030 MHz Receiver
In the frequency band of this survey, commercially available devices were able to provide adequate receiver sensitivity and bandwidth. Low-noise amplifiers were followed by conventional amplification and band-equalizing stages placed on the feed platform. The receiver passband was defined by the cut-off frequency of the feed at ${\sim}350$ MHz, and by a low-pass filter, with 3 dB frequency ${\sim}1030$ MHz, placed at the end of the feed-platform amplifier chain. The amplified outputs were transported to the control building by a radio frequency (RF)-over-fiber (RFoF) device. Signals were further amplified to 20-20 dBm, the specified input level to the analog-to-digital converters. The calibration signal was provided by a solid-state noise source heated to a constant temperature near $30\arcdeg$ C. [^2] Outside nighttime temperatures did not exceed $30\arcdegC$ during the survey, and it is simpler to only heat the noise source rather than also implementing cooling in warmer weather.
II.3 The Digital Spectrometer
@@
Since the key data product from this survey is a Faraday depth cube, the requirements for the spectral channelization are primarily driven by bandwidth depolarization effects (Raja, 2014; Schnitzeler and Lee, 2015; Pratley and Johnston-Hollitt, 2020), which could, in principle, be met by a spectrometer that analyzes incoming signals to a frequency resolution of ${\sim}1$ MHz (see Section VI.2 for details). Faraday synthesis on such data produces a reliable Faraday depth cube (and we ultimately bin the channels to 1 MHz in the map-making stage). Also, digitization to four bits would satisfactorily limit any additional noise from the quantization process. However, in the real world, our observing frequencies carry heavy communications traffic, and we need much finer digitization in order to have the dynamic range to detect the astronomical signal in the presence of the much stronger radio frequency interference (RFI) signals. Channels and samples spoiled by RFI must be excised from the data, and to minimize data loss it is useful to have a sampling that is much finer than required by the demands of Faraday synthesis. Fortunately, modern digital electronics permits the design of spectrometers with many bits and many narrow frequency channels.
@@
Two of the 137.5 MHz-wide bands lie below the cutoff frequency of the feed and therefore do not contain useful data. The number of spectrometer channels used is about 16,300. Inter-channel isolation (between 41.6 kHz channels) is 70 dB, and the passband ripple is 0.1 dB. Aliasing at the 1030 MHz roll-off of the low-pass filter is ${\sim}40$ dB. The four correlated outputs produced by the spectrometer are the total power from the left- and right-circular channels, RR* and LL*, and the linear polarization products, RL* and LR*.
II.4 Telescope Control System
Survey observations were made as azimuth scans at two fixed elevations (Section III.1). With its azimuth-elevation configuration, the DRAO-15 telescope has the capability to scan quickly in azimuth (up to 180 deg/minute) at a fixed elevation. The azimuth axis uses two drives to prevent backlash. [^3] Discrepancy in the motion due to the clearance between the gear teeth. The elevation drive has a maximum speed of 60 deg/minute, allowing the telescope to change between the two survey elevations ($49\arcdeg$ and $20\arcdeg$; Section III.1) in less than 4 minutes (slow-down and settling time included), permitting efficient scheduling of scans at the two elevations used for the survey within an observing night.
The low-level control software was written in C++, while the observing schedule for the survey (including azimuth scans and calibration scans) consisted of a set of python scripts. For the cross-scans for the pointing model (Section II.5) and the raster scans of bright calibrators (Section III.2.2) python scripts provided the telescope with a series of pointings in right ascension (RA) and declination, updated every 0.5 seconds, tracing out the pattern of the scan by commanding the telescope to ‘chase’ the changing target pointing.
Figure 4: Right ascension (top) and declination (bottom) offsets between the expected and observed pointing positions as a function of azimuth. These measurements are derived from Gaussian fits to the calibration raster scans of Cyg A, Cas A, Tau A, and Vir A (Section III.2.2) produced over the course of the survey.
II.5 Pointing Model
@@
Each observation consisted of a ‘cross-scan’: one $8\arcdeg$-long scan in RA and one $8\arcdeg$-long scan in declination across the expected source position. Peak positions of 1D Gaussian functions fitted to each scan in the LL* correlation product at frequencies above 900 MHz determined the measured positions. The six model parameters fitted to the data with TPoint were: azimuth offset, elevation offset, non-perpendicularity between the azimuth and elevation axes, elevation-dependent azimuth shift, azimuth-axis east-west misalignment and azimuth-axis north-south misalignment. After applying the corrections we re-observed a smaller subset of sources in order to assess the new pointing accuracy. The combined corrections produced a substantial improvement, reducing the systematic offsets between the expected and observed source positions by $\sim 80\%$ in both elevation and azimuth, and reducing the rms errors by $6\%$ in elevation and by $50\%$ in azimuth. The overall residual uncertainty in pointing is $\sim 5\%$ of the beamwidth, making this a sufficiently accurate correction for mapping large-scale structures.
@@
We applied the pointing model at the beginning of the survey and no further corrections were made throughout. However the daily calibration raster scans using the four brightest sources (Cyg A, Cas A, Tau A, and Vir A; Section III.2.2) provided continuous checks on the pointing accuracy. We show the RA and declination offsets in Figure 3 as a function of frequency based on an averaged set of Cyg A raster scans, and in Figure 4 at 820 MHz as a function of telescope azimuth from the full set of raster scans. While there is a slight residual pointing accuracy dependence on both frequency and azimuth, the pointing accuracy remains within 5% of the beamwidth over the course of the survey, indicating sufficient pointing stability.
II.6 System and Noise-Source Temperature
@@
Separate from the calibration measurements and procedures that we ultimately applied to the survey data (described in Section III.2), we measured the noise temperature of the receiver and the noise equivalent of the injected calibration signal prior to installing the receiver on the telescope. This was accomplished using the DRAO Hot-Cold Test Facility (HCTF; Hovey et al., 2018). The HCTF is a large metal funnel, with its wide mouth open at the top. The floor of the funnel is a flat metal plate, 1 m square, and the top opening is 3 m square, 3 m above the floor. The feed and receiver were placed in it, pointing upwards. Noise measurements were put on an absolute scale by exposing the feed and receiver to two terminations at known temperatures. A hot termination consists of a sliding roof lined with microwave absorber, and a cold termination was provided by sliding the roof open, exposing feed and receiver to the sky. We measured the physical temperature of the hot absorber directly, and derived the brightness temperature of the sky from the PyGDSM sky model (Price, 2016; Zheng et al., 2017). Measurements were made with laboratory spectrum analyzers with a frequency resolution of 3 MHz.
@@
Figure 5: The receiver noise temperature in kelvins (top) and the kelvin equivalent of the calibration noise source (bottom) as determined from the HCTF measurements. Results are shown for both right- and left-hand circular polarization (RCP; black and LCP; gray).
@@
The receiver temperature and calibration noise-source temperature are shown in Figure 5. The receiver temperature varies smoothly between $\sim 80$ K at the high and low ends of the band and $\sim 55$ K at 600 MHz. The calibration noise-source temperature fluctuates between $\sim 15$ and $\sim 80$ K. The large-scale, sinusoidal fluctuations in the noise-source temperature are likely caused by reflections between the noise-source coupler (Section III.2.1) and the metal plate to which it was mounted. We estimate the accuracy of these measurements as ${\pm}3$ K above 500 MHz. Accuracy below 500 MHz is limited by the size of the ${\sim}45$  cm-long pyramidal cones of the absorber that line the sliding roof. To investigate the effects of aliasing into the band we repeated the HCTF measurements out to 4 GHz, revealing that a negligible 0.05 to 0.3 K would be added to the system temperature by aliasing.
We calculate the expected noise in the final 1 MHz channel widths used for published data cubes using the radiometer equation applied to the receiver temperature values in Figure 5 and approximate sky brightness temperatures from the PyGDSM sky model (Price, 2016; Zheng et al., 2017). An integration time of $2\times 0.6=1.2$ s [^5] The effective integration time is twice the spectrometer integration time since each pointing is sampled twice in the scan strategy (Section III.1). and 1 MHz channel-width yields an expected noise between 55 and 100 mK.
II.7 Aperture Efficiency
As described in Section IV.8, we achieved the conversion to brightness temperature (K) by calculating the main beam solid angle, as a function of frequency, from the measured beamwidths, assuming that the main beam is Gaussian. Nevertheless, we carried out an absolute calibration of the telescope, with the aim of improving our understanding of the instrument. We compared the antenna temperature produced by Cyg A with absolute noise standards, resistors at the temperature of liquid nitrogen, at ambient temperature, and in a temperature controlled oven at approximately $100^{\circ}$ C. We made this measurement using the RR* beam only. In the top panel of Figure 6 we show the measured aperture efficiency, ${\eta}_{A}$, compared with that computed using the General Reflector antenna Analysis Software Package by TICRA (GRASP) [^6] https://www.ticra.com/software/grasp/ electromagnetic simulator. The excellent agreement, generally within 1%, gives us great confidence in the GRASP computations of the radiation pattern of the antenna, which will be useful for future improvements to the ground emission calculations described in Section IV.4.
III Observations
@@ III.1 Observing Strategy
We used the observing technique developed by Carretti et al. (2019), scanning in azimuth at fixed elevation. The timing was selected such that scans adjacent to each other on the sky provided close to uniform spatial coverage, and the length of the scans was chosen such that scans would also cross one another, allowing for the possibility of basket-weaving to reconcile their base levels (see Section IV.6). A scan consists of all the data collected for one sweep in azimuth with the telescope moving either clockwise or counterclockwise over $360^{\circ}$ at $20^{\circ}$ minute-1, starting and ending at azimuth ${a}={0^{\circ}}$ (north). Scans were made at two elevations, ${e}={49.3^{\circ}}$, the elevation of the North Celestial Pole (NCP) at DRAO, and at ${e}={20^{\circ}}$, just above the lower elevation limit for the telescope. We denote these as ${e}_{\text{high}}$ and ${e}_{\text{low}}$ scans respectively. The ${e}_{\text{high}}$ scans covered the declination range $11\arcdeg\leq\delta\leq 90\arcdeg$ and ${e}_{\text{low}}$ scans covered $20\arcdeg\leq\delta\leq 60\arcdeg$ (Figure 7). In the map-making stage we split each scan into ‘east’ ($0\arcdeg\leq{a}\leq 180\arcdeg$) and ‘west’ ($180\arcdeg\leq{a}\leq\ 360\arcdeg$) halves (Figure 8) which produced the intersecting scan patterns (Figure 9). We produced separate maps from the ${e}_{\text{high}}$ and ${e}_{\text{low}}$ datasets. We then combined these using a weighted average in the overlap region $11\arcdeg\leq\delta\leq\ 60\arcdeg$, with the ${e}_{\text{high}}$ (${e}_{\text{low}}$) map weights changing linearly from 0 to 1 (1 to 0) across $\delta=11\arcdeg$ to $\delta=60\arcdeg$, thereby producing the final maps covering $20\arcdeg\leq\delta\leq 90\arcdeg$.
@@
[…]
@@
In the survey planning process we generated a ‘library’ of scans, following Wolleben et al. (2021), with each scan having a unique 4-digit scan ID. This consisted of 1440 ${e}_{\text{low}}$ and 1440 ${e}_{\text{high}}$ scans. For each phase of the survey we generated a daily scan schedule, drawing from the scans with sidereal start times occurring after sunset and sidereal end times occurring before sunrise. These schedules were generated with two objectives: (i) to achieve survey efficiency, and (ii) to add some randomness to the scan sequence. These objectives were, obviously, in conflict, so we adopted a reasonable compromise, aiming to avoid fixed patterns in the observing sequence. Sets of both ${e}_{\text{high}}$ and ${e}_{\text{low}}$ scans were completed on each night. After each survey night, if a scan was deemed unsuccessful (e.g. a result of excessive RFI or equipment malfunction) the scan ID was returned to the master scan library to be rescheduled on a subsequent night. In total, 1401 of the 1440 ${e}_{\text{high}}$ and 1406 of the 1440 ${e}_{\text{low}}$ scans were completed for inclusion in the mapping stage. We made efforts to reschedule failed scans, but were in some cases limited by a lack of available nighttime coverage of the required sidereal observing times. In the processing steps we also discarded a small number of scans from individual frequency channels based on an rms filtering threshold (Section IV.3).
III.2 Calibration Observations